OpenStudy (anonymous):

Evaluate when x = 2, y = 7 and z = 6: 4(x+y)-7 A. 20 B. 10 C. 30 D. 25 @alibaby

OpenStudy (anonymous):

so im supposed to explain it or give u the answer

OpenStudy (anonymous):

explai i just need to know how to do it thats all

OpenStudy (anonymous):

ok hold on

OpenStudy (ryuga):

https://mathway.com/ is a good site

OpenStudy (anonymous):

What do we get when we replace x with 2, y with 7 and z with 6 in this expression? \(4(x+y)-7\)

OpenStudy (anonymous):

u do () first but how do u do the 4 in this equation

OpenStudy (anonymous):

Expression Calculator evaluates an expression in a given context. Context of evaluation is specified by a comma separated list of equations. Both symbolical and numerical computations are supported. Calculations with complex numbers are also supported. If the intent is for the Expression Calculator to compute a numberical answer, then context should contain sufficient number of equations to resolve all variables in your expression. Expression 4*(x+y)-7 = 4*y+4*x-7 Copy to clipboard 4y+4x−7 Expression x=2[y=7*andz=6]

OpenStudy (anonymous):

do u see mine @de

OpenStudy (anonymous):

@DERRICKCURRY

OpenStudy (anonymous):

yea it helped meah

OpenStudy (anonymous):

You didn't need to distribute it, that was unnecessary. \[4(x+y)-7\]\[4(2+7)-7\]\[4\times9-7\]Can you simplify that?

OpenStudy (anonymous):

yea is equals 30

OpenStudy (anonymous):

the 7 in the equation is supose to be z

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yep, well done.

OpenStudy (anonymous):

dont forget to close @DERRICKCURRY